OpenStudy (anonymous):

HELP whats the simplified form of the square root of 150b^3??????

OpenStudy (anonymous):

would 1837.1 be the answer, wat does the "b" mean...?

OpenStudy (anonymous):

@Izabelle.Esther it has to be simplified not the actual answer it has to be either 6b square root of 5b 5b square rot of 6 5b square root of 5 or 5b square root of 6

OpenStudy (anonymous):

OHH now i get wat ur doin, SORRY MY BAD>>>>

zepdrix (zepdrix):

\[\large \sqrt{150b^3} \qquad = \qquad \sqrt{25\cdot6\cdot b^2\cdot b} \qquad = \qquad \sqrt{25b^2}\cdot\sqrt{6b}\]

OpenStudy (mathstudent55):